OpenStudy (anonymous):

square root 153 minus square root 17 plus square root 68 simplify if possible

OpenStudy (anonymous):

= 3sqrt17 - sqrt17 + 2sqrt17 = 4 sqrt17

OpenStudy (mathmate):

sqrt(153)-sqrt(17)+sqrt(68) =sqrt(3^2*17)-sqrt(17)+sqrt(2^2*17) =3sqrt(17)-sqrt(17)+2sqrt(17) =4sqrt(17)

OpenStudy (anonymous):

Answer= 4sqrt(17)

OpenStudy (anonymous):

\[4\sqrt{17}\]